OpenStudy (anonymous):

\[\sqrt{x+13}=x+1 ?\]

OpenStudy (tkhunny):

Please note, first, that \(x+13\ge 0\). You tell me why.

OpenStudy (anonymous):

Becasue 13 and anything added to that is greater than 0

OpenStudy (anonymous):

?

OpenStudy (tkhunny):

Okay, that's not wrong, but why do we care and why did I mention it?

OpenStudy (anonymous):

Dont we need to figure a number to make both sides satisfied?

OpenStudy (tkhunny):

Yes we do. If we pick a number less than -13, the left-hand side doesn't exist. It is important to know this. If the result of our algebra is x = -25, we can just throw it out. That will not work. Okay, now square both sides.

OpenStudy (anonymous):

\[x^{2}+26=x^{2}+2 ?\]

OpenStudy (tkhunny):

\(\left(\sqrt{x+13}\right) = x+13\) I really do not know how you managed that previous response. \((x+1)^{2} = x^{2} + 2x + 1\) I think you may have forgotten your algebra. \(x+13 = x^{2} + 2x + 1\)

OpenStudy (anonymous):

Thanks, and yes thats why i am here...for HELP!

OpenStudy (tkhunny):

You should be learning the solution of radical equations. You shoudl already have as tools the squaring of radicals and the squaring of binomials. you are supposed to bring tools with you. Learning everything all over again is not a good plan. So, what solution have you managed?